# Blade: A Derivative-free Bayesian Inversion Method using Diffusion Priors

### v1

@@ 3.4 Theoretical analysis
In this section, we analyze the non-asymptotic behavior of Blade through the lens of its continuous-time and large particle limit for the ease of understanding. In practice, Blade incurs two bias terms: $\epsilon_{\mathrm{model}}$ from the statistical linearization and $\epsilon_{\mathrm{score}}$ from the learned diffusion prior. By extending existing proof techniques from Wu et al. (2024); Vempala & Wibisono (2019); Sun et al. (2024), our analysis quantifies how these errors together with the number of iterations $K$ affects the deviation from the reference process. The technical definitions and notations are collected in Appendix A.1.
@@
Given any $\rho>0$, consider the oracle split-Gibbs algorithm that alternates between the likelihood step defined in Eq. (6) and the prior step defined in Eq. (2) where each step is implemented perfectly without approximations. If the particle distribution is not a Dirac measure, then $\pi^{XZ}$ is a stationary distribution. Furthermore, if the particle covariance $C_{t}$ is positive definite, $\pi^{XZ}$ is the unique stationary distribution.
@@
Let $T_{k}=k(t^{\dagger}+t^{*}),k=0,\dots,K$, $\lambda^{*}=\inf_{t\in\cup_{k}[T_{k},T_{k}+t^{\dagger}]}\lambda^{*}_{t}$, and $\delta=\inf_{t\in[0,t^{*}]}\delta(t)$ where $\delta(t)$ is the diffusion term defined in Eq. (14). We denote by $\epsilon_{\mathrm{score}}$ the score approximation error of the diffusion model defined in Assumption 1, and $\epsilon_{\mathrm{model}}$ the forward model derivative approximation error defined in Assumption 2. Assuming that $D_{\mathrm{KL}}(\pi^{X}||\mu_{0})<+\infty$ and Assumption 3 holds, for $K$ split Gibbs iterations, we have
@@
The complete proofs of Theorems  1 and 2 are deferred to Appendix A.2.
@@
Theorem 1 show that, when every sub-step is executed exactly, $\pi^{XZ}$ is a stationary distribution. Theorem 2 accounts for the effect of the two generally unavoidable approximations: the statistical linearization of the forward model and the learned diffusion prior. Eq. (12) indicates that the time-average Fisher divergence between the approximate process and reference process decays at an $O(1/K)$ rate up to a weighted sum of two approximation errors. While this bound is structurally similar to that of PnPDM (Wu et al., 2024), there are two key distinctions. First, our method considers fundamentally different likelihood step dynamics (with statistical linearization and covariance preconditioning); consequently, the resulting error term depends on additional factors such as covariance matrix eigenvalues and the linearization error. Second, unlike the analysis in Wu et al. (2024) which assumes exact likelihood steps, our analysis incorporates the effects of model error $\epsilon_{\mathrm{model}}$ and a finite time horizon $t^{\dagger}$.
@@ 4 Experiments
Inverse problems are ill-posed, so what matters in practice is often not a single point estimate but calibrated posterior samples. We therefore evaluate Blade through probabilistic verification of its posterior samples. Sec. 4.1 considers fully controlled settings with analytic ground-truth posteriors, enabling direct distributional checks. Sec. 4.2 turns to a challenging high-dimensional problem based on the Navier-Stokes equation where the ground-truth posterior is unknown but probabilistic verification methods are available. Sec. 4.3 studies the effect of Blade’s hyperparameters and its robustness to diffusion prior choice. Sec. D reports results on popular point-estimate benchmarks as complementary evidence of Blade’s breadth.
@@ 4.1 Gaussian and Gaussian mixture
To enable direct distributional check, we consider two cases where the exact posteriors can be derived: (i) a linear Gaussian posterior and (ii) a two‐component non-isotropic Gaussian mixture posterior. For both we derive the analytic posterior and draw “ground‐truth” samples, then compare to those produced by different algorithms. We compare Blade against the existing derivative-free methods including DPG (Tang et al., 2024), SCG (Huang et al., 2024), EnKG (Zheng et al., 2025a), and EKS (Garbuno-Inigo et al., 2020a). Fig. 1 shows scatter plots of the joint samples alongside marginal density estimates. In the linear Gaussian test (Fig. 1, top row), Blade and EKS, as sampling methods, are able to capture the posterior mean and variance. In contrast, optimization-based algorithms like SCG, DPG, and EnKG operate by minimizing a surrogate objective and thus return point estimates, failing to capture posterior spread/uncertainty. In the multimodal non-isotropic Gaussian mixture case (Fig. 1, 2nd row), Blade recovers both modes and their relative weights while EKS yields an over-dispersed Gaussian approximation. This is because the particles of EKS share a single potential function while Blade assign individual potentials to particles, enabling multi-modal posterior sampling (as mentioned in Sec. 3.1). The detailed setup, derivations of the analytic posteriors, and additional evaluation metrics, are in Appendix C.1, with further quantitative results in Appendix 3.
@@ 4.2 Navier-Stokes equation
To test Blade in a high-dimensional and challenging setting that is closer to the real-world problems, we consider the problem of recovering the initial vorticity field in the two-dimensional Navier–Stokes equations from partial, noisy observations taken at a later time. We pick this problem because the dynamics is highly nonlinear and mirrors many practical challenges in science and engineering including weather data assimilation (White, 2000), geophysics (Liu & Gurnis, 2008), and fluid reconstruction (Elsinga et al., 2006). More importantly, this setting has standard probabilistic verifications available to test the quality of posterior samples.
Table 1: Comparison on the Navier-Stokes inverse problem. The primary probabilistic metrics are CRPS and SSR. Rel L2 error (relative L2 error) is deterministic, included as a complementary metric. - indicates either that probabilistic metrics are inapplicable (deterministic method) or that it is too costly to generate enough samples from the algorithm for reliable calculation of probabilistic metrics.
@@
Problem setup. We take the Navier-Stokes problem formulation in InverseBench (Zheng et al., 2025b) where a non-trivial distribution is considered. The initial vorticity $\bm{x}$$^{*}$ in resolution $128\times 128$ is evolved forward with a numerical solver, then subsampled and corrupted with Gaussian noise of standard deviation $\sigma_{\mathrm{noise}}=0,1,2$. The observation $\bm{y}$} thus constitute a partial, noisy snapshot of the flow field. We use the publicly released dataset and pretrained diffusion prior from InverseBench. All the experiments are conducted on single GH200 GPU. Details are in Appendix C.2.
Baselines. We compare our algorithm against two classes of methods. The first class is the methods that only requires diffusion prior trained on unpaired data including DPG (Tang et al., 2024), SCG (Huang et al., 2024), EnKG (Zheng et al., 2025a), EKI (Iglesias et al., 2013), EKS (Garbuno-Inigo et al., 2020a) (initialized from diffusion prior). The second class is provided as reference points, which requires additional training on paired data, including conditional diffusion model (CDM) and an end-to-end UNet. The conditional diffusion learns the posterior distribution through conditional score matching. The UNet directly learns to predict ground truth from the observation. For each noise regime we retrain both the conditional diffusion model and the U-Net from scratch, using the same training configuration. The details are in Appendix C.2.3.
Evaluation metrics. For comprehensive evaluation, we consider three different metrics in the literature (Zheng et al., 2025b; Rasp et al., 2024) to assess the performance from both probabilistic and deterministic perspectives: continuous ranked probability score (CRPS), spread-skill ratio (SSR), and relative L2 error (Rel L2 error). CRPS is a proper scoring rule that rewards both sharp predictive distribution and well-calibrated predictions whereas SSR diagnoses calibration only. An SSR near one is desirable but must be interpreted in conjunction with the other error metrics. Formal definitions and implementation details of these metrics are in Appendix C.2.2.
@@ 4.3 Ablation studies
Data scaling. We evaluate how the performance of Blade depends on the pre-trained diffusion prior. We retrain the diffusion prior with progressively larger data subsets (ranging from $10\times 2^{7}$ to $10\times 2^{11}$). We compare the resulting performance to that of the end-to-end UNet and CDM, both retrained for each subset and each noise level. As shown in Fig. 10, Blade surpasses the baselines by a significant margin. In contrast, the UNet improves only modestly with data size. These results indicate that Blade is sample-efficient, can efficiently exploit extra prior data when available yet maintaining descent performance even in low-data regime.
@@ 5 Conclusion
Acknowledgments
H.Z. is supported by the PIMCO and Amazon AI4science fellowship. Z.W. is supported by the Amazon AI4Science fellowship.
@@ A.2 Proofs
Remark 3.
Assumption 1 coincides with the usual bounded score-matching error condition that underpins convergence results (Lee et al., 2022; Chen et al., 2022; Wu et al., 2024). Assumption 2 characterizes the L2 accuracy of the linear proxy. Assumption 3 is the weakest assumption to bound the weighted Fisher divergence in our analysis. Two common sufficient (but not necessary) scenarios are: (1) $C_{t}$ is full-rank (2) $\mu_{t}$ is absolutely continuous with respect to $\tilde{\mu}_{t}$ and both log densities are continuously differentiable.
@@
This result has been proved in various forms in the literature (Ma et al., 2015; Garbuno-Inigo et al., 2020a), we provide a simple proof of our use case for ease of understanding. Suppose $\mu_{t}({\bm{z}})$ is the probability density of $\bm{z}$} at time $t$. For the ease of notation, we ignore the particle index $j$ in ${\bm{z}}_{t}$. Let $\Phi({\bm{z}})=f({\bm{z}};{\bm{y}})+\frac{1}{2\rho^{2}}\|{\bm{z}}-{\bm{x}}^{(j)}\|_{2}^{2}$. The corresponding Fokker-Planck equation for Eq. (6) reads
@@
If the particle distribution is not Dirac, $C_{t}\neq 0$ due to Lemma 2.1 in Garbuno-Inigo et al. (2020a). Therefore,
@@
showing that $\pi^{Z|X={\bm{x}}^{(j)}}({\bm{z}})$ is a stationary distribution of the dynamics of Eq. (6). Further, if $C_{t}$ is positive definite, it ensures the irreducibility and strong Feller property, and the stationary distribution is unique (Roberts & Tweedie, 1996; Ma et al., 2015). ∎
@@
This is a generalization of Lemma A.4 in Wu et al. (2024) to the general matrix-valued diffusion term. Intuitively, the condition that $b_{t}-\tilde{b}_{t}$ belongs to the range of $C(t)$ means that the two drift terms may only differ along the directions that are actually driven by noise. In the context of our proof below, this is always satisfied because the drift terms are either preconditioned with $C(t)$ or $C(t)$ is full-rank.
@@
This proof is largely adapted from Nüsken & Reich (2019); Garbuno-Inigo et al. (2020b) which applies to more general scenarios. We tailor and simplify the proof for our use case for the ease of understanding.
@@ Appendix B Practical implementation
In this section, we detail the practical implementation of each component of the proposed method, Blade. We also present ablation studies for each design choice to elucidate their individual effects. All the ablation studies are conducted on on a small subset of the Navier-Stokes inverse problem’s test set.
@@ B.1 Likelihood step
As Theorem 2 indicates, the initialization of Blade is quite flexible, provided the initial distribution maintains a finite KL divergence from the target distribution. Our empirical evaluation considered two initialization strategies: Gaussian and diffusion prior (DM) initialization. For the Navier-Stokes inverse problem, As shown in Fig. 12, the empirical performance difference between two initializations is not substantial. In contrast, for image restoration tasks, we observed that Gaussian initialization gets better results. This is because the target distribution being more closely approximated by a Gaussian due to the large initial coupling strength. Conversely, DM initialization tends to produce a natural image distribution, which is known to often exhibit an unbounded KL divergence from the Gaussian (Arjovsky & Bottou, 2017). Therefore, we use Gaussian initialization for the image restoration tasks.
@@
During the likelihood step, we employ a resampling strategy to ensure that the particles are at the correct noise level $\rho$. Resampling is a commonly used method that has been shown to help improve the performance of algorithms such as DAPS Zhang et al. (2025), DiffPIR Zhu et al. (2023), and ReSample Song et al. (2023). Specifically, we define the following resampling strategy:

### v2

@@ 3.4 Theoretical analysis
In this section, we analyze the behavior of Blade through the lens of its continuous-time and large particle limit for the ease of understanding. In practice, Blade incurs two bias terms: $\epsilon_{\mathrm{model}}$ from the statistical linearization and $\epsilon_{\mathrm{score}}$ from the learned diffusion prior. By extending existing proof techniques from wu2024principled; vempala2019rapid; sun2024provable, our analysis quantifies how these errors together with the number of iterations $K$ affects the deviation from the reference process. Technical definitions and notations are collected in Appendix A.1.
@@
Given any $\rho>0$, consider the oracle split-Gibbs algorithm that alternates between the likelihood step defined in Eq. (6) and the prior step defined in Eq. (2) where each step is implemented perfectly without approximations. If the particle distribution is not a Dirac measure, then $\pi^{XZ}$ is a stationary distribution. Furthermore, if the preconditioner $C_{t}$ is positive definite, $\pi^{XZ}$ is the unique stationary distribution.
We defer the proof to Appendix A.2.
Remark 2.
Theorem 1 shows that, when every sub-step is exact, $\pi^{XZ}$ is the stationary distribution under the standard argument with positive definite preconditioner. Note that Theorem 1 holds for any positive definite preconditioner and the covariance matrix is just one special case. The preconditioner does not change the stationary distribution but rather specifies the geometry of the dynamics (See Kalman-Wasserstein gradient flow in garbuno2020interacting). This observation will be used in Theorem 2 to compare two processes within a common metric structure.
The next theorem establishes convergence and stability of Blade under the two practical approximation errors. The prior score approximation error arising from the diffusion model is defined in Assumption 1, which coincides with the usual bounded score-matching error condition that underpins convergence results (lee2022convergence; chen2022sampling; wu2024principled). The statistical linearization error $\epsilon_{\mathrm{model}}$ defined in Assumption 2 characterizes the L2 accuracy of the statistical linearization. It is uniformly bounded under the standard regularity assumptions on forward model and uniformly bounded covariance matrix. We also make Assumption 3, which is the weakest condition to bound the weighted Fisher divergence in our analysis. Two common sufficient (but not necessary) scenarios are: (1) $C_{t}$ is full-rank (2) $\mu_{t}$ is absolutely continuous with respect to $\tilde{\mu}_{t}$ and both log densities are continuously differentiable.
@@
Let $T_{k}=k(t^{\dagger}+t^{*}),k=0,\dots,K$, $\lambda^{*}=\inf_{t\in\cup_{k}[T_{k},T_{k}+t^{\dagger}]}\lambda^{*}_{t}$, and $\delta=\inf_{t\in[0,t^{*}]}\delta(t)$ where $\delta(t)$ is the diffusion term defined in Eq. (14). We assume both $\lambda^{*}$ and $\delta$ are strictly positive. We denote by $\epsilon_{\mathrm{score}}$ the score approximation error of the diffusion model defined in Assumption 1, and $\epsilon_{\mathrm{model}}$ the forward model derivative approximation error defined in Assumption 2. Assuming that $D_{\mathrm{KL}}(\pi^{X}||\mu_{0})<+\infty$ and Assumption 3 holds, for $K$ split Gibbs iterations, we have
@@
The proof of is deferred to Appendix A.2.
@@
Theorem 2 provides two main insights. The first is the convergence to the reference process. The time-average Fisher divergence between the approximate process and reference process decays at an $O(1/K)$ rate up to a weighted sum of approximation errors. Second, the algorithm remains stable under the statistical linearization error and the prior score approximation error and these terms do not vanish with more iterations as expected.
Interpretation of the reference process. The reference process stays at the target stationary distribution if the likelihood step is exact regardless of the chosen preconditioner $C_{t}$. The role of preconditioner $C_{t}$ here is to specify the geometry with respect to which we measure approximation error as mentioned in Remark 2. Using $C_{t}$ from the approximate process to construct the reference process does not alter the target stationary distribution; it simply allows us to compare the two dynamics within a common metric structure. Hence, the convergence to the reference process indicates the convergence to the desired stationary distribution
Comparison to prior work While the bound in Eq. 12 is structurally similar to that of PnPDM (wu2024principled), our analysis strictly generalizes the prior results to a more realistic and technically challenging regime. First, our method considers a different and more complex likelihood step dynamics with interacting particles and state-dependent preconditioner, whereas wu2024principled considers the standard Langevin dynamics. Second, our setting is more realistic by considering finite time execution of the dynamics as well as the forward model error. Third, Theorem 2 explicitly shows how the algorithm remains stable under forward model error and how the finite time horizon affect convergence, which are not addressed in the prior work.
@@ 4 Experiments
Inverse problems are ill-posed, so an ensemble of posterior samples with calibrated uncertainty is often more desirable than a single point estimate. We therefore evaluate Blade through probabilistic verification of its posterior samples. Sec. 4.1 considers fully controlled settings with analytic ground-truth posteriors, enabling direct distributional checks. Sec. 4.2 turns to a challenging high-dimensional problem based on the Navier-Stokes equation where the ground-truth posterior is unknown but probabilistic verification methods are available. Sec. 4.3 studies the effect of Blade’s hyperparameters and its robustness to diffusion prior choice. Sec. D reports results on popular point-estimate benchmarks as complementary evidence of Blade’s breadth.
@@ 4.1 Gaussian and Gaussian mixture
To enable direct distributional check, we consider two cases where the exact posteriors can be derived: (i) a linear Gaussian posterior and (ii) a two‐component non-isotropic Gaussian mixture posterior. For both we derive the analytic posterior and draw “ground‐truth” samples, then compare to those produced by different algorithms. We compare Blade against the existing derivative-free methods including DPG (tang2024solving), SCG (huang2024symbolic), EnKG (zheng2024ensemble), and EKS (garbuno2020interacting). Fig. 1 shows scatter plots of the joint samples alongside marginal density estimates. In the linear Gaussian test (Fig. 1, top row), Blade and EKS, as sampling methods, are able to capture the posterior mean and variance. In contrast, optimization-based algorithms like SCG, DPG, and EnKG operate by minimizing a surrogate objective and thus return point estimates, failing to capture posterior spread/uncertainty. In the multimodal non-isotropic Gaussian mixture case (Fig. 1, 2nd row), Blade recovers both modes and their relative weights while EKS yields an over-dispersed Gaussian approximation. This is because the particles of EKS share a single potential function while Blade assign individual potentials to particles, enabling multi-modal posterior sampling (as mentioned in Sec. 3.1). The detailed setup, derivations of the analytic posteriors, and additional evaluation metrics, are in Appendix C.1, with further quantitative results in Appendix 3.
@@ 4.2 Navier-Stokes equation
To test Blade in a high-dimensional and challenging setting that is closer to the real-world problems, we consider the problem of recovering the initial vorticity field in the two-dimensional Navier–Stokes equations from partial, noisy observations taken at a later time. We pick this problem because the dynamics is highly nonlinear and mirrors many practical challenges in science and engineering including weather data assimilation (white2000ifs), geophysics (liu2008simultaneous), and fluid reconstruction (elsinga2006tomographic). More importantly, this setting has standard probabilistic verifications available to test the quality of posterior samples.
Table 1: Comparison on the Navier-Stokes inverse problem. The primary probabilistic metrics are CRPS and SSR. Rel L2 error (relative L2 error) is deterministic, included as a complementary metric. - indicates either that probabilistic metrics are inapplicable (deterministic method) or that it is too costly to generate enough samples from the algorithm for reliable calculation of probabilistic metrics. CDM-CA: conditional diffusion model with cross attention. CDM-Cat: conditional diffusion model with channel concatenation.
@@
Problem setup. We take the Navier-Stokes problem formulation in InverseBench (zhenginversebench) where a non-trivial distribution is considered. The initial vorticity $\bm{x}$$^{*}$ in resolution $128\times 128$ is evolved forward with a numerical solver, then subsampled and corrupted with Gaussian noise of standard deviation $\sigma_{\mathrm{noise}}=0,1,2$. The observation $\bm{y}$} thus constitute a partial, noisy snapshot of the flow field. We use the publicly released dataset and pretrained diffusion prior from InverseBench. All the experiments are conducted on single GH200 GPU. Details are in Appendix C.2.
Baselines. We compare our algorithm against two classes of methods. The first class is the methods that only requires diffusion prior trained on unpaired data including DPG (tang2024solving), SCG (huang2024symbolic), EnKG (zheng2024ensemble), EKI (iglesias2013ensemble), EKS (garbuno2020interacting) (initialized from diffusion prior), EKS W/ DP (EKS with diffusion prior), localized EKS w/ DP (reich2021fokker; wagner2022ensemble), and FKD (singhal2025general; zhao2025conditional). The second class is provided as reference points, which requires additional training on paired data, including conditional diffusion model (CDM) and an end-to-end U-Net. The conditional diffusion learns the posterior distribution through conditional score matching. The U-Net directly learns to predict ground truth from the observation. For each noise regime we retrain both the conditional diffusion model and the U-Net from scratch, using the same training configuration. The details are in Appendix C.2.3.
Evaluation metrics. For comprehensive evaluation, we consider three different metrics in the literature (zhenginversebench; rasp2024weatherbench) to assess the performance from both probabilistic and deterministic perspectives: continuous ranked probability score (CRPS), spread-skill ratio (SSR), and relative L2 error (Rel L2 error). CRPS is a proper scoring rule that rewards both sharp predictive distribution and well-calibrated predictions whereas SSR diagnoses calibration only. An SSR near one is desirable but must be interpreted in conjunction with the other error metrics. Formal definitions and implementation details of these metrics are in Appendix C.2.2.
@@ 4.3 Ablation studies
Data scaling. We evaluate how the performance of Blade depends on the pre-trained diffusion prior. We retrain the diffusion prior with progressively larger data subsets (ranging from $10\times 2^{7}$ to $10\times 2^{11}$). We compare the resulting performance to that of the end-to-end U-Net and CDM, both retrained for each subset and each noise level. As shown in Fig. 10, Blade surpasses the baselines by a significant margin. In contrast, the U-Net improves only modestly with data size. These results indicate that Blade is sample-efficient, can efficiently exploit extra prior data when available yet maintaining decent performance even in low-data regime.
@@ A.2 Proofs
This result has been proved in various forms in the literature (ma2015complete; garbuno2020interacting), we provide a simple proof of our use case for ease of understanding. Suppose $\mu_{t}({\bm{z}})$ is the probability density of $\bm{z}$} at time $t$. For the ease of notation, we ignore the particle index $j$ in ${\bm{z}}_{t}$. Let $\Phi({\bm{z}})=f({\bm{z}};{\bm{y}})+\frac{1}{2\rho^{2}}\|{\bm{z}}-{\bm{x}}^{(j)}\|_{2}^{2}$. The corresponding Fokker-Planck equation for Eq. (6) reads
@@
If the particle distribution is not Dirac, $C_{t}\neq 0$ due to Lemma 2.1 in garbuno2020interacting. Therefore,
@@
showing that $\pi^{Z|X={\bm{x}}^{(j)}}({\bm{z}})$ is a stationary distribution of the dynamics of Eq. (6). Further, if $C_{t}$ is positive definite, it ensures the irreducibility and strong Feller property, and the stationary distribution is unique (roberts1996exponential; ma2015complete). ∎
@@
This is a generalization of Lemma A.4 in wu2024principled to the general matrix-valued diffusion term. Intuitively, the condition that $b_{t}-\tilde{b}_{t}$ belongs to the range of $C(t)$ means that the two drift terms may only differ along the directions that are actually driven by noise. In the context of our proof below, this is always satisfied because the drift terms are either preconditioned with $C(t)$ or $C(t)$ is full-rank.
@@
This proof is largely adapted from nusken2019note; garbuno2020affine which applies to more general scenarios. We tailor and simplify the proof for our use case for the ease of understanding.
@@ Appendix B Practical implementation
In this section, we detail the practical implementation of each component of the proposed method, Blade. We also present ablation studies for each design choice to elucidate their individual effects. All the ablation studies are conducted on a small subset of the Navier-Stokes inverse problem’s test set.
@@ B.1 Likelihood step
As Theorem 2 indicates, the initialization of Blade is quite flexible, provided the initial distribution maintains a finite KL divergence from the target distribution. Our empirical evaluation considered two initialization strategies: Gaussian and diffusion prior (DM) initialization. For the Navier-Stokes inverse problem, as shown in Fig. 12, the empirical performance difference between two initializations is not substantial. In contrast, for image restoration tasks, we observed that Gaussian initialization gets better results. This is because the target distribution being more closely approximated by a Gaussian due to the large initial coupling strength. Conversely, DM initialization tends to produce a natural image distribution, which is known to often exhibit an unbounded KL divergence from the Gaussian (arjovsky2017principledmethodstraininggenerative). Therefore, we use Gaussian initialization for the image restoration tasks.
@@
During the likelihood step, we employ a resampling strategy to ensure that the particles are at the correct noise level $\rho$. Resampling is a commonly used method that has been shown to help improve the performance of algorithms such as DAPS zhang2024improving, DiffPIR zhu2023denoising, and ReSample song2023solving. Specifically, we define the following resampling strategy: